Paul and Maureen, two graduates at University, produce customized personal
computer systems at night. They put together generic personal computers, which have various
special features for students, and local businesses. They have recently received their biggest
order to date from the EM department for 36 customized personal computers. Paul and
Maureen assembled the first unit as a trial and found that it took them 18 hours of direct labor.
To determine if they can fill the order in the time allotted, they want to apply the learning curve
effect to determine how much time the 9th, 18th, and 36th units will require to assemble. Based
on past experience, they believe their learning curve is 80%
Answer #1

Answer:

Time required for the 9th unit is computed using learning curve formula: (tn =t1nb)

t9 = (18)(9)ln(0.8)/ln2

= (18)(9)-0.322

=(18)/(9)0.322

= 18*0.493

=8.874 hrs

the times required for 18th and 26th units are computed similarly

t18 = (18)(18)ln(0.8)/ln2

= 18*0.394

=7.092 hrs

t36 = (18)(36)ln(0.8)/ln2

= 18*0.315

=5.67 hrs

Thank you.